OpenStudy (anonymous):

Which of the following is a polynomial with roots -2, -2, and 2 ? x3 - 4x2 + 4x - 16 x3 + 4x2 + 4x + 16 x3 + 2x2 + 4x + 8 x3 - 2x2 + 4x - 8

jhonyy9 (jhonyy9):

x^2(x+2)+4(x+2)=(x^2 +4)(x+2) x^2 +4=0 x_1,_2=+/- 2 x+2=0 x=-2 so than will be 3rd equation right

OpenStudy (anonymous):

yes thank you